40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 22 02:43:28 2011, kiwamu, Update, General, in-vac work 
    Reply  Mon Aug 22 21:32:10 2011, kiwamu, Update, General, more in-vac work : AS clipping fixed and OSEM/oplev adjustment 
       Reply  Tue Aug 23 06:49:24 2011, Anamaria, Update, General, more in-vac work : AS clipping fixed and OSEM/oplev adjustment 
          Reply  Tue Aug 23 09:40:37 2011, kiwamu, Update, General, Re: AS clipping fixed 
Message ID: 5285     Entry time: Tue Aug 23 09:40:37 2011     In reply to: 5284
Author: kiwamu 
Type: Update 
Category: General 
Subject: Re: AS clipping fixed 

Indeed it was suspenseful.

We tried finding where the clipping happened, but we couldn't find any obvious clippings.

So we checked centering of the beams on all the optics associated with the AS path, starting from BS, SR3,... to the AS optical bench.

And during the work some of them were recentered.

At the end we found no clipping. To make sure we tested the available range (no clipping range) by exciting the angular motion of BS with AWG (f ~ 1Hz, a ~ 1000).

The beam looked successfully coming out at the most of the angular oscillation point.

Quote from #5284

Where was the AS clipping?! Ah, the suspense...

 

ELOG V3.1.3-